OpenStudy (anonymous):

Find the first 5 terms of the sequence: an = n2 - 2n. Enter your answers separated by commas; example: 1, -5, 10, -15, 20

OpenStudy (anonymous):

I get 1, -1, 1, -1, 1... I feel like that is wrong though

OpenStudy (anonymous):

JK I got -1, 3, 3, 3, 3... Still seems wrong

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

JK both of those are wrong...

OpenStudy (anonymous):

I got it, it is -1, 0, 3, 8, 15

OpenStudy (anonymous):

ok, let check n=1, then \(a_1 = 1^2 -2*1 = -1\) n =2, then \(a_2 = 2^2 -2*2 = 0\) n =3 , then \(a_3 = 3^2 -2*3 = 3\)

OpenStudy (anonymous):

ok, I can stop now, hihihi

OpenStudy (anonymous):

haha thank you though haha

OpenStudy (here_to_help15):

1) It will depend on whether you start with a(0) or a(1). If you start with a(0), we will have: 0, -1, 0, 3, 8 If you start with a(1), we will have: -1, 0, 3, 8, 15